OpenStudy (anonymous):

find the value of the logarithmic function without using a calculator log[27]^9 a. 3 b. 3/2 c. 2/3 d. 1/3

OpenStudy (ybarrap):

log(27)^9, using log base 3 = 9*log(27) = 9*log(3^3) = 9*3*log(3) = 27 * log(3) = 27 * 1 = 27 Ans. 27, using the convenient base of 3.